9. Value added to a firm Financial statements reflect only book values of the data that analysts use to evaluate a companysYou have taken a job as an entry-level analyst, and your boss has asked you to find the expected value of Adams Furnitures s
Answer #1

EVA = Net Operating profit After tax – Invested Capital*Cost of capital

= (80,000,000-52,000,000)(1-40%) – 83,500,000*7.5%

= $10,537,500

Hence, value added during the year = $10,537,500

Value of Stock = Next Expected Dividend/(Required Rate of Return – Growth Rate)

= 2.75(1+3%)/(11%-3%)

= $35.40625

i.e. $35.41

Value = 2.75(1+3%)/(14%-3%)

= $25.75

Fall in value by $9.66